# Disentangling quantum autoencoder

### v1

@@ IV Training
Next, we discuss how to train the DQAE such that it can disentangle a given distribution of quantum states $W$. The training will be performed in an unsupervised manner, i.e. the quantum data has no labels. We sample $L$ training states $|\psi^{(\ell)}\rangle\in W$ as the training set $S=\{|\psi^{(\ell)}\rangle\}_{\ell=1}^{L}$ from $W$. The training states are then used to train the DQAE. The DQAE is trained by a parameterized unitary $U(\boldsymbol{\theta})$ via $U(\boldsymbol{\theta})|\psi^{(\ell)}\rangle$, where $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{M})$ is a $M$-dimensional parameter vector. We minimize a cost function $C_{\text{train}}(\boldsymbol{\theta})$ by varying the parameter $\boldsymbol{\theta}$. The cost function is given by the average purity of the reduced single-qubit density matrix
$C_{\text{train}}(\boldsymbol{\theta})$ $=1-\frac{1}{LN}\sum_{\ell=1}^{L}\sum_{k=1}^{N}E_{k}(U(\boldsymbol{\theta})|\psi^{(\ell)}\rangle)\,$ (13)
$E_{k}(|\psi\rangle$ $=\text{tr}(\text{tr}_{\bar{k}}(|\psi\rangle\langle\psi|)^{2})\,,$ (14)
where $\text{tr}_{\bar{k}}(.)$ is the partial trace over all qubits except the $k$th qubit. We have $C_{\text{train}}(\boldsymbol{\theta})=0$ only when $U(\boldsymbol{\theta})|\psi^{(\ell)}\rangle$ is a pure single-qubit product state for all $\ell$, else $C_{\text{train}}(\boldsymbol{\theta})>0$.
@@
where $\sigma_{k}^{\alpha}$ with $\alpha\in\{x,y,z\}$ are the Pauli operators acting on the $k$th qubit. The Pauli expectation can be easily estimated on a quantum computers by measuring the respective eigenbasis of the Pauli operators, which requires measuring in $3$ different basis ($x$,$y$, and $z$).
As alternative approach, we can use Bell measurements on two copies of the state $|\psi\rangle$. We define the Bell state $|\Psi^{-}\rangle=\frac{1}{\sqrt{2}}(|01\rangle-|10\rangle)$, the $2$-qubit identity matrix $I_{2}$ and the projector $\Pi_{k}=I_{2}^{\otimes k-1}|\Psi^{-}\rangle\langle\Psi^{-}|I_{2}^{\otimes N-k}$ onto the Bell state of the $k$th qubit pair. It can be shown that [39, 40, 41]
@@
This can be achieved by measuring only in the Bell basis, as the projectors for all $k$ commute.
We now consider a general parameterized quantum circuits of $d$ layers of the form $|\psi(\boldsymbol{\theta})\rangle=\prod_{j=1}^{d}V_{j}(\boldsymbol{\theta}_{n})T_{j}|0\rangle$ with fixed entangling unitary $T_{j}$, parameter vector $\boldsymbol{\theta}$ and parameterized rotations $V_{j}(\theta_{j})=e^{-i\boldsymbol{\theta}_{j}/2\sigma_{j}}$ generated by some Pauli string $\sigma_{j}$. For such circuits, gradients $\partial_{j}E_{k}(|\psi(\boldsymbol{\theta})\rangle)$ can be measured efficiently with the shift-rule [42]. The shift-rule can be directly adapted to purity measurements [41]
@@ V Overparameterization and generalization
Generalization can be similarly evaluated using the DQFIM. An overparameterized model generalizes (assuming no noise) when the rank of the DQFIM does not increase further upon increasing the number $L$ of training states. Note that for $L=1$, the DQFIM is identical to the well known quantum Fisher information metric [44].
Let use first introduce the data state for training set $S_{L}=\{|\psi^{(\ell)}\rangle\}_{\ell=1}^{L}$
@@
For a given $L$ training states, the maximal rank of the DQFIM is given as
$R_{L}\equiv\max_{M\geq M_{\text{c}}(L),\boldsymbol{\theta}}\text{rank}(\mathcal{Q}(\rho_{L},U(\boldsymbol{\theta})))\,.$ (20)
where $M_{\text{c}}(L)$ is the critical number of circuit parameters where the model overparameterizes. For $M\geq M_{\text{c}}(L)$, training is likely to converge to the global minima. Note that $M_{\text{c}}(L)$ depends on the number $L$ of training data, and in general the circuit parameters needed to overparameterize increase with $L$.
For generalization (i.e. low test error), one has to supply a critical number of training data $L\geq L_{\text{c}}$ during training
$R_{\infty}=R_{L_{\text{c}}}\equiv\max_{L\geq L_{\text{c}}}R_{L}(\rho_{L},U)\,,$ (21)
such that $R_{L}$ does not increase further upon increase of $L$ (for $M\geq M_{\text{c}}(L)$). The maximal possible $R_{L}$ is the maximal rank of the DQFIM $R_{L_{\text{c}}}\equiv R_{\infty}$. The critical number of training states needed to generalize can be approximated as [28]
@@
Thus, the rank of the DQFIM allows us to determine how many circuit parameters $M_{\text{c}}$ and training states $L_{\text{c}}$ are needed to generalize, which we will later on apply for the DQAE.
@@
Figure 5: Parameterized unitaries for disentangler unitary $U(\boldsymbol{\theta})$, which are composed of many layers of the following building blocks: a) Generic hardware-efficient ansatz $U_{\text{HE}}$ of parameterized $y$ and $z$ rotations and fixed CNOT gates arranged in nearest-neighbor fashion. b) Fermionic ansatz $U_{\text{fermion}}$ c) Transverse-field Ising inspired ansatz $U_{\text{Ising}}$.
@@ VI Ansatz circuits
We study different types of ansatz circuits as shown in Fig. 5. In Fig. 5a, we have a generic hardware efficient ansatz, constructed from CNOT gates in a 1D nearest-neighbor fashion $W$ combined with layer of parameterized rotations generated from Pauli operators $\sigma^{x}_{j}$, $\sigma^{z}_{j}$ acting on qubit $j$, which is written as
@@ VII DQAE with stabilizer states
Stabilizer states and Clifford circuits are one of the fundamental building blocks of quantum information and computing [46]. Due to their simpler structure, they have an efficient algorithm to compute many properties, and even their dynamics can be efficiently learned from quantum states [47]. They are constructed from the Clifford gate set: A combination of CNOT, Hadamard and S-gates, which can be efficiently simulated classically [46].
Here, we propose Algorithm 1 to learn to disentangle stabilizer states, which is inspired from Ref. [48] and the Metropolis algorithm. The goal is to build a disentangling unitary, starting from the identity operator. The idea of the algorithm is to randomly apply either a random single-qubit Clifford unitaries $\mathcal{C}_{1}$ or CNOT gate on randomly chosen qubits. This gate is accepted with probability $p=\text{min}(1,\exp(-(C_{\text{p}}-C)/T(b)))$, where $T(b)$ is a temperature function which monotonously decreases with each step $b$, $C_{\text{p}}$ is the cost function with the added gate, and $C$ without. These steps are repeated until the cost function reaches zero. We empirically find that this algorithm nearly always converges, and we find only a weak dependence on performance depending on the chosen $T(b)$.
For stabilizer input data, the quantum part of the algorithm can be efficiently simulated on classical computers [49]. Note that a classical representation of stabilizer states can be efficiently learned on quantum computers [50]. Thus, we can run this algorithm by first learning a classical representation of the training set, then training the disentangling unitary on a classical computer, where the quantum part is simulated classically.
@@ VIII.1 Data generation
To generate training and test data, we first draw a fixed unitary $V$ from the respective studied ansatz as shown in Fig. 5. Then, for the training data we prepare $L$ random single-qubit product states $|\psi^{(\ell)}\rangle=\bigotimes_{j}|\phi_{j}^{(\ell)}\rangle$, which gives us the training set
@@
which is used to train the circuit. The test set is drawn similarly, which is not used during training and only for testing the trained ansatz. Our choice of $S_{L}$ ensures that there exists an optimal solution with low test and training error. However, this solution may be difficult to find during training when the number of training data $L$ is low, or the circuit does not have enough circuit parameters $M$.
@@
Figure 6: We show the variance of the gradient for training the DQAE against qubit number $N$. The gradient is computed at random parameters $\boldsymbol{\theta}$ averaged over 10 random instances. We show a generic ansatz $U_{\text{HE}}$, fermionic ansatz $U_{\text{fermion}}$ and Ising model ansatz $U_{\text{ising}}$. Dashed line is fit with $\text{var}(\nabla C_{\text{train}})=cN^{-\gamma}$, where we find $\gamma\approx 3.8$ for $U_{\text{fermion}}$ and $U_{\text{ising}}$.
@@ VIII.2 Trainability
Now, we study the trainability of the DQAE, i.e. whether the gradients for training are sufficiently large [25]. We show the variance of the gradient $\text{var}(\nabla C_{\text{train}})$ against $N$ in Fig. 6 for different ansatz types as detailed in Fig. 5. We find that for generic hardware efficient ansatz, the variance decays exponentially which indicates a barren plateau [22]. For $U_{\text{ising}}$ and $U_{\text{fermion}}$ we find the gradient decays only polynomially, indicating the absence of barren plateaus for these ansatz unitaries, and the circuit can be trained even when the number of qubits is scaled up.
@@
Figure 7: Training error $C_{\text{train}}$ (solid line) and test error $C_{\text{test}}$ (dashed) for circuit ansatz $U_{\text{HE}}$ as shown in Fig. 5a. We show error as function of number of training epochs for different number of training states $L$ for $M=400$ circuit parameter in the overparameterized regime. We have critical number of training states $L_{\text{c}}=16$ and $N=4$.
Refer to captionaRefer to caption Refer to captionbRefer to caption Refer to captioncRefer to caption
Figure 8: DQAE for disentangling a generic circuit ansatz $U_{\text{HE}}$ as shown in Fig. 5a. We show the final a) training loss $C_{\text{train}}$, b) test loss $C_{\text{test}}$ and c) number of iterations $E$ needed for training to converge. We show $N=4$ qubits, for different number of training states $L$ and circuit parameters $M$. Dashed lines are thresholds for overparameterization via Eq. (20) and generalization via Eq. (22)
VIII.3 Generic ansatz
Next, we study the generic ansatz, shown in Fig. 5a, in more detail. First, we show a typical training trajectory for the YZ-ansatz in Fig. 7 for different number of training states $L$. We find three regimes in the overparameterized regime $M\gg M_{\text{c}}$: For $L\ll L_{\text{c}}$, $C_{\text{train}}$ converges fast, but the test error remains high. For $L\gg L_{\text{c}}$, dataset is overcomplete and both $C_{\text{train}}$ and $C_{\text{test}}$ converge fast. For $L=15\approx L_{\text{c}}$, we observe that $C_{\text{train}}$ takes a long time to converge, with $C_{\text{test}}$ becoming smaller as well over many epochs.
Next, we study the training and test cost for the generic ansatz for different number of circuit parameters $M$ and training data size $L$. find that the training cost in Fig. 8a becomes small when the circuit is about to overparameterized, i.e. $M>M_{\text{c}}(L)$, where the threshold from the DQFIM is indicated as the horizontal-aligned dashed line. Furthermore, as shown in Fig. 8b, the test error decreases significantly once the model generalizes, which occurs when the number of training states exceeds $L_{\text{c}}\approx 2R_{\infty}/R_{1}$, as indicated by the vertical dashed line. We find that around the overparameterization and generalization threshold, the number of epochs needed to train substantially increases as shown in Fig. 8c. In general, for a generic circuit ansatz, both $L_{\text{c}}$ and $M_{\text{c}}(L)$ scale exponentially with the number of qubits $N$. This has been analytically calculated in Ref. [28], where one finds $L_{\text{c}}=2^{N}$ and $M_{\text{c}}(L)=\text{min}(2^{N+1}L-L^{2}-1,4^{N}-1)$.

### v2

@@ IV Training
Next, we discuss how to train the DQAE such that it can disentangle a given distribution of quantum states $W$. The training will be performed in an unsupervised manner, i.e. the quantum data has no labels. We sample $L$ training states $|\psi^{(\ell)}\rangle\in W$ as the training set $S_{L}=\{|\psi^{(\ell)}\rangle\}_{\ell=1}^{L}$ from $W$. The training states are then used to train the DQAE. The DQAE is trained by a parameterized unitary $U(\boldsymbol{\theta})$ via $U(\boldsymbol{\theta})|\psi^{(\ell)}\rangle$, where $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{M})$ is a $M$-dimensional parameter vector. We minimize a cost function $C_{\text{train}}(\boldsymbol{\theta})$ by varying the parameter $\boldsymbol{\theta}$. The cost function is given by the average purity of the reduced single-qubit density matrix
$C_{\text{train}}(\boldsymbol{\theta})$ $=1-\frac{1}{LN}\sum_{\ell=1}^{L}\sum_{k=1}^{N}E_{k}(U(\boldsymbol{\theta})|\psi^{(\ell)}\rangle)\,.$ (15)
We have $C_{\text{train}}(\boldsymbol{\theta})=0$ only when $U(\boldsymbol{\theta})|\psi^{(\ell)}\rangle$ is a pure single-qubit product state for all $\ell$, else $C_{\text{train}}(\boldsymbol{\theta})>0$.
@@
where $\sigma_{k}^{\alpha}$ with $\alpha\in\{x,y,z\}$ are the Pauli operators acting on the $k$th qubit. The Pauli expectation can be easily estimated on a quantum computer by measuring the respective eigenbasis of the Pauli operators, which requires measuring in $3$ different bases ($x$,$y$, and $z$).
As an alternative approach, we can use Bell measurements on two copies of the state $|\psi\rangle$. We define the Bell state $|\Psi^{-}\rangle=\frac{1}{\sqrt{2}}(|01\rangle-|10\rangle)$, the $2$-qubit identity matrix $I_{2}$ and the projector $\Pi_{k}=I_{2}^{\otimes k-1}|\Psi^{-}\rangle\langle\Psi^{-}|I_{2}^{\otimes N-k}$ onto the Bell state of the $k$th qubit pair. It can be shown that [39, 40, 41]
@@
This can be achieved by measuring only in the Bell basis, as the projectors for all $k$ commute. The purity $E_{k}$ can be estimated with $O(1/\epsilon^{2})$ samples within $\epsilon$ additive accuracy via either Pauli or Bell-basis measurements, which can be easily shown using Hoeffding’s inequality. As $E_{k}$ for different $k$ commute, all $E_{k}$ can be measured in parallel, thus the overall number of samples needed to estimate $C_{\text{train}}$ is independent of $N$.
We now consider a general parameterized quantum circuit of $d$ layers of the form $|\psi(\boldsymbol{\theta})\rangle=\prod_{j=1}^{d}V_{j}(\theta_{j})T_{j}|0\rangle$ with fixed entangling unitary $T_{j}$, parameter vector $\boldsymbol{\theta}$ and parameterized rotations $V_{j}(\theta_{j})=e^{-i\theta_{j}/2\sigma_{j}}$ generated by some Pauli string $\sigma_{j}$. For such circuits, gradients $\partial_{j}E_{k}(|\psi(\boldsymbol{\theta})\rangle)$ can be measured efficiently with the shift-rule [42]. The shift-rule can be directly adapted to purity measurements [41]
@@ V Overparameterization and generalization
Generalization can be similarly evaluated using the DQFIM. An overparameterized model generalizes (assuming no noise) when the rank of the DQFIM does not increase further upon increasing the number $L$ of training states. Note that for $L=1$, the DQFIM is identical to the well-known quantum Fisher information metric [44].
Let us first define the data state for a given training set $S_{L}=\{|\psi^{(\ell)}\rangle\}_{\ell=1}^{L}$ [28]
@@
For a given set of $L$ training states, the maximal rank of the DQFIM is given as [25, 28]
$R_{L}\equiv\max_{M\geq M_{\text{c}}(L)}\max_{\boldsymbol{\theta}}\text{rank}(\mathcal{Q}(\rho_{L},U(\boldsymbol{\theta})))\,.$ (21)
We say the circuit is overparameterized for a given $L$ when increasing $M$ does not increase $R_{L}(M)$. The smallest such value defines the critical number of circuit parameters $M_{\text{c}}(L)$, i.e., [25, 28]
$R_{L}(M)=R_{L}(M_{\text{c}}(L))\quad\text{for all }M\geq M_{\text{c}}(L)\,.$ (22)
For overparameterized $M\geq M_{\text{c}}(L)$, training is likely to converge to the global minima. Note that $M_{\text{c}}(L)$ depends on the number $L$ of training data, and the circuit parameters needed to overparameterize increase with $L$.
For generalization (i.e. low test error), one has to also supply a critical number of training data $L\geq L_{\text{c}}$ during training. The maximal rank is defined as [28]
$R_{\infty}=R_{L_{\text{c}}}\equiv\max_{L\geq L_{\text{c}}}R_{L}(\rho_{L},U)\,$,
The critical number of training states $L_{\text{c}}$ is given by the smallest $L$ such that the DQFIM rank saturates
$R_{L}=R_{\infty}\quad\text{for all }L\geq L_{\text{c}}\,.$ (23)
where $L_{\text{c}}$ is the minimum number of training states that are needed in order to achieve low test error. To simplify calculation, it can be approximated by [28]
@@
Thus, the rank of the DQFIM allows us to determine how many circuit parameters $M_{\text{c}}$ and training states $L_{\text{c}}$ are needed to generalize, which we will later apply for the DQAE. Note that deep circuits usually concentrate, i.e. random instances of the circuit assume the average value of the ensemble. Thus, determining $R_{L}$ numerically is straightforward without optimization: One chooses a few random $\theta$ and picks the largest found rank of the DQFIM, which, due to the concentration, is likely to be the maximal rank. Note that $\text{rank}(\mathcal{Q}$ and $R_{L}$ increases monotonously with $M$ and $L$ (in fact usually nearly linear as seen in Fig.2 of Ref. [28]), thus determining $M_{\text{c}}(L)$ and $L_{\text{c}}$ can be done straightforwardly. In many cases, the rank of the DQFIM is known analytically, such as for hardware-efficient circuits or circuits with a known dynamical Lie algebra [28].
@@
Figure 5: Parameterized unitaries for disentangler unitary $U(\boldsymbol{\theta})$, which are composed of many layers of the following building blocks: a) Hardware-efficient ansatz $U_{\text{HE}}$ of parameterized $y$ and $z$ rotations and fixed CNOT gates arranged in nearest-neighbor fashion. b) Fermionic ansatz $U_{\text{fermion}}$ c) Transverse-field Ising inspired ansatz $U_{\text{Ising}}$.
@@ VI Ansatz circuits
We study different types of ansatz circuits as shown in Fig. 5. In Fig. 5a, we have a hardware-efficient ansatz, constructed from CNOT gates in a 1D nearest-neighbour fashion $W$ combined with layers of parameterized rotations generated from Pauli operators $\sigma^{x}_{j}$, $\sigma^{z}_{j}$ acting on qubit $j$, which is written as
@@ VII DQAE with stabilizer states
Stabilizer states and Clifford circuits are one of the fundamental building blocks of quantum information and computing [46]. Due to their simpler structure, they have an efficient algorithm to compute many properties, and even their dynamics can be efficiently learned from quantum states [47]. They are constructed from the Clifford gate set: a combination of CNOT, Hadamard, and S-gates, which can be efficiently simulated classically [46].
Here, we propose Algorithm 2 to learn to disentangle stabilizer states, which is inspired by Ref. [48] and the Metropolis algorithm. The goal is to build a disentangling unitary, starting from the identity operator. The idea of the algorithm is to randomly apply either a random single-qubit Clifford unitary $\mathcal{C}_{1}$ or CNOT gate on randomly chosen qubits. This gate is accepted with probability $p=\text{min}(1,\exp(-(C_{\text{p}}-C)/T(b)))$, where $T(b)$ is a temperature function which monotonically decreases with each step $b$, $C_{\text{p}}$ is the cost function with the added gate, and $C$ without. These steps are repeated until the cost function reaches zero. We empirically find that this algorithm nearly always converges, with only a weak dependence on performance depending on the chosen $T(b)$.
For stabilizer input data, the quantum part of the algorithm can be efficiently simulated on classical computers [49]. Note that a classical representation of stabilizer states can be efficiently learned on quantum computers [50] using $O(N)$ copies of the state. Thus, we can run our Metropolis algorithm efficiently as follows: First, we learn the stabilizer states of the training set, which can be done using Ref. [50] from $O(N)$ measurements. In fact, any stabilizer state can be represented using its generators of the stabilizing Pauli operators, which takes $O(N)$ classical memory. Then, we train the disentangling unitary on a classical computer, where the quantum part is simulated classically using the Tableau formalism for Clifford unitaries [49].
@@ VIII.1 Data generation
To generate training and test data, we first draw a fixed unitary $V$ from the respective studied ansatz, as shown in Fig. 5. Then, for the training data we prepare $L$ random single-qubit product states $|\psi^{(\ell)}\rangle=\bigotimes_{j}|\phi_{j}^{(\ell)}\rangle$, which gives us the training set
@@
which we use to train our circuit. The test set is drawn similarly, which is not used during training and only for testing the trained ansatz. Our choice of $S_{L}$ ensures that there always exists a DQAE with zero test and training error. However, this solution may be difficult to find during training for the optimization algorithm: For example, convergence to the global minimum is unlikely when the circuit is not overparameterized, i.e. $M<M_{\text{c}}$. Further, the test error usually remains high when the number of training data is low, i.e. $L<L_{\text{c}}$.
@@
Figure 6: We show the variance of the gradient for training the DQAE against qubit number NN. The gradient is computed for random parameters $\boldsymbol{\theta}$ averaged over 10 random instances. We show a hardware-efficient ansatz $U_{\text{HE}}$, fermionic ansatz $U_{\text{fermion}}$ and Ising model ansatz $U_{\text{ising}}$. Dashed line is fit with $\langle\text{var}(\partial_{j}C_{\text{train}})\rangle=cN^{-\gamma}$, where we find $\gamma=3.80\pm 0.46$ for $U_{\text{fermion}}$, and $\gamma=3.87\pm 0.5$ for $U_{\text{ising}}$.
@@ VIII.2 Trainability
Now, we study the trainability of the DQAE, i.e. whether the gradients for training are sufficiently large [25]. We show the variance of the gradient $\langle\text{var}(\partial_{j}C_{\text{train}})\rangle$ against $N$ in Fig. 6 for different ansatz types as detailed in Fig. 5. Here, we average the variance over all circuit parameters. We find that for hardware-efficient ansatze, the variance decays exponentially, which indicates a barren plateau [22]. For $U_{\text{ising}}$ and $U_{\text{fermion}}$, we find the gradient decays only polynomially, indicating the absence of barren plateaus for these ansatz unitaries, and the circuit can be trained even when the number of qubits is scaled up.
@@
Figure 7: Training error $C_{\text{train}}$ (solid line) and test error $C_{\text{test}}$ (dashed) for circuit ansatz $U_{\text{HE}}$ as shown in Fig. 5a. We show the error as a function of the number of training epochs for different numbers of training states $L$ for $M=400$ circuit parameters, where we are in the overparameterized regime for $N=4$. We have critical number of training states $L_{\text{c}}=2^{N}=16$ as derived in Ref. [28].
Refer to captionRefer to caption
Figure 8: DQAE for disentangling a hardware-efficient ansatz $U_{\text{HE}}$ as shown in Fig. 5a. We show the final a) logarithm of training loss $C_{\text{train}}$, b) logarithm of test loss $C_{\text{test}}$ and c) number of epochs needed for training to converge to error $10^{-3}$. We show $N=4$ qubits, for different numbers of training states $L$ and circuit parameters $M$. Dashed lines are thresholds for overparameterization via Eq. (21) and generalization via Eq. (24), which are known analytical as shown in Eq. (28).
VIII.3 Hardware-efficient ansatz
Next, we study a hardware-efficient ansatz, shown in Fig. 5a, in more detail. First, we show a typical training trajectory for the YZ-ansatz in Fig. 7 for different numbers of training states $L$. We find three regimes, where we consider the overparameterized limit $M\gg M_{\text{c}}$: For $L\ll L_{\text{c}}$, $C_{\text{train}}$ converges fast, but the test error remains high. For $L\gg L_{\text{c}}$, dataset is overcomplete and both $C_{\text{train}}$ and $C_{\text{test}}$ converge fast to zero. For $L=15\approx L_{\text{c}}$ close to the critical dataset size, we observe that $C_{\text{train}}$ takes a long time to converge, with $C_{\text{test}}$ converging slowly over many epochs.
Next, we study the training and test cost for the hardware-efficient ansatz for different numbers of circuit parameters $M$ and training data size $L$. We find that the training cost in Fig. 8a becomes small when the circuit is about to become overparameterized, i.e. $M>M_{\text{c}}(L)$, where the threshold from the DQFIM is indicated as the horizontally aligned dashed line. Furthermore, as shown in Fig. 8b, the test error decreases significantly once the model generalizes, which occurs when the number of training states exceeds $L_{\text{c}}\approx 2R_{\infty}/R_{1}$, as indicated by the vertical dashed line. We find that around the overparameterization and generalization threshold, the number of epochs needed to train substantially increases, as shown in Fig. 8c. In general, for a hardware-efficient ansatz, both $L_{\text{c}}$ and $M_{\text{c}}(L)$ scale exponentially with the number of qubits $N$. In fact, in this case $R_{L}$ can be computed analytically, yielding
$M_{\text{c}}(L)=\text{min}(2^{N+1}L-L^{2}-1,4^{N}-1)$ (28)
and $L_{\text{c}}=2^{N}$ [28].